# Single-Domain Generalized Object Detection by Balancing Domain Diversity and Invariance

### v1

@@ Abstract
Single-domain generalization for object detection (S-DGOD) aims to transfer knowledge from a single source domain to unseen target domains. In recent years, many models have focused primarily on achieving feature invariance to enhance robustness. However, due to the inherent diversity across domains, an excessive emphasis on invariance can cause the model to overlook the actual differences between images. This overemphasis may complicate the training process and lead to a loss of valuable information. To address this issue, we propose the Diversity Invariance Detection Model (DIDM), which focuses on the balance between the diversity of domain-specific and invariance cross domains. Recognizing that domain diversity introduces variations in domain-specific features, we introduce a Diversity Learning Module (DLM). The DLM is designed to preserve the diversity of domain-specific information with proposed feature diversity loss while limiting the category semantics in the features. In addition, to maintain domain invariance, we incorporate a Weighted Aligning Module (WAM), which aligns features without compromising feature diversity. We conducted our model on five distinct datasets, which have illustrated the superior performance and effectiveness of the proposed model.
1 Introduction
@@
In recent years, deep learning has led to significant improvements in the performance of detection models. However, traditional detectors often rely on the assumption of Independent and Identically Distributed (i.i.d.) data. This assumption becomes a limitation when the model encounters a domain shift, leading to a significant drop in performance and restricting the applicability of these models in real-world scenarios.
Figure 1: A brief overview of DIDM is presented. The DLM is designed for the domain-specific feature to preserve their diversity and suppress the semantic information. The WAM is employed to align features without compromising the feature diversity with loss weight.
To address the domain shift problem, domain generalization (DG) has been developed to transfer knowledge from the source domains to the unseen target domain [10, 25, 49, 59]. These methods enhance model robustness, making it more robust to more application scenarios. Recently, single-domain generalization [26, 41, 42, 58, 60] has gained significant attention as it reduces data requirements by training the model on a single domain, which is a much more challenging task compared to the traditional DG.
Most existing single-domain generalization models aim to reduce domain bias by learning domain-invariant features, which often restrict the model and prevent the capture of domain-specific information. This approach requires the model to extract similar, or even identical, features from images across different domains. Considering significant domain differences, such a constraint is counterintuitive and can limit effective feature extraction, potentially complicating the training process. To address this, we proposed that a domain generalization (DG) model should not only consider the learning of domain invariance but also the diversity of the domain-specific information. The learning of domain invariance ensures cross-domain robustness, while the diversity of domain-specific features enhances feature learning and model training.
@@
Based on the above discussion, specifically, in this paper, we introduce the Diversity Invariance Domain Generalization Detection Model (DIDM), designed to not only learn the domain invariance but also keep the diversity domain-specific. To manage these two feature types, we incorporate the Diversity Learning Module (DLM) to capture domain-specific features and the Weighted Aligning Module (WAM) to focus on domain-invariant features, as illustrated in Fig. 1.
The Diversity Learning Module (DLM) is designed to preserve the domain diversity and limit the semantic limitation. Firstly, to eliminate the semantic information among the domain-specific features, we apply an Entropy Maximization loss to reduce category-level semantics. However, focusing solely on constraining domain-specific information may limit the model’s ability to learn diverse features. To address this, we introduce a feature diversity loss (FD) to reinforce the diversity of domain-specific information. Finally, the DLM offers three key advantages: first, by reducing category semantics in domain-specific features, it supports domain-invariant training; second, it preserves feature diversity across domains, enhancing the training phase of the model. Additionally, by encouraging feature diversity with FD loss, the module compels the model to capture richer semantic information from images, resulting in more semantically nuanced features.
In addition to domain-specific feature learning with DLM, domain invariance learning may also influence the feature diversity in the Domain Generalization (DG) model. Thus, we introduce the Weighted Aligning Module (WAM) to alleviate diversity loss with feature alignment. The proposed WAM is designed with two main objectives: First, by feature alignment across domains, it enables the model to learn domain-invariant features for the detection task; Second, by incorporating loss weights, the model prevents the overemphasizes the feature alignment. When the feature differences do not influence the detection, the model down-weight the alignment to preserve the domain diversity. By combining both the DLM and WAM, our model achieves effective domain generalization on the detection task. In conclusion, the contribution of the proposed DIDM is presented as follows:
  • We introduce the Diversity Learning Module (DLM) to preserve the feature diversity while limiting the semantic information within domain-specific features. The feature diversity loss is proposed to enhance the feature diversity.
  • The Weighted Aligning Module (WAM) is designed to align features without influencing the feature diversity. The loss weight is proposed to prevent an overemphasis on domain-specific features.
  • Experiments were conducted under five distinct weather conditions, effectively showcasing the robustness of our methodology.
@@
2 Related Work
Domain Generalization (DG). Researchers have been working extensively in single domain generalization and have developed numerous methods to enhance it [10, 25, 28, 42, 49]. These methods can generally be grouped into three main categories: First, the data argumentation approach [19, 20, 21, 34] increases the diversity of source domain data by applying various data manipulation techniques. For instance, Somavarapu et al. [34] introduce a straightforward image stylization transform to generate diverse samples, exploring variability across sources. Similarly, Wang et al. [40] employ adversarial training to create varied input images, further enhancing generalization. Second, some kinds of DG methods focus on the presentation learning of domain invariance. Shao et al. [35] use multi-adversarial discriminative training to extract both shared and distinctive feature representations across multiple source domains. Last, Various learning strategies [2, 8, 30, 37, 44, 47] are employed to improve model generalization. For example, Zhao et al. [47] apply a meta-learning approach that simulates the model’s adaptation to new, unseen domains during training, thereby boosting its adaptability in unfamiliar environments. These approaches illustrate significant advancements in domain generalization, highlighting different techniques designed to improve models’ performance on previously unseen domains.
Object Detection. To mitigate the impact of domain bias on model performance, researchers have conducted numerous experiments and proposed various methods in object detection. One prominent approach is Domain Adaptive Object Detection (DAOD) [6, 11, 18, 27, 29, 50], which enhances the model’s robustness in new domains by simultaneously training on both source and target domains. For instance, a common strategy involves minimizing the distance between global and local features, as outlined in [3, 4, 11, 36]. However, these methods require access to the target domain during model training, which imposes limitations on practical applications. Therefore, researchers have proposed Domain Generalized Object Detection (DGOD) [61, 62]. For instance, Chang et al. [61] enhance the model’s generalization ability by decoupling depth estimation from dynamic perspective enhancement. However, the effectiveness of DGOD largely depends on the number of accessible source domains, and collecting multiple source domains can be costly. Consequently, researchers have introduced the more challenging Single Domain Generalized Object Detection (S-DGOD) [60, 12, 58, 31, 7], which is generally categorized into feature normalization and invariant-based methods. Firstly, feature normalization methods such as IBN-Net, proposed by Pan et al. [31], enable the network to adjust its normalization strategy according to different tasks and datasets by combining Instance Normalization (IN) and Batch Normalization (BN). IterNorm, introduced by Huang et al. [14], avoids feature decomposition through a Newton iteration method, thereby improving the efficiency of the normalization process. Secondly, invariant-based approaches, such as UFR proposed by Liu et al. [25], enhance the model’s generalization performance by eliminating prototype bias and attentional bias. Wu et al. [41] propose a cyclic disentanglement self-distillation approach specifically for single-domain generalization in object detection, which enhances feature disentanglement. These approaches demonstrate significant progress in the field of target detection and contribute to the rapid advancement of the discipline.
3 Methodology
In this section, we introduce the proposed Diversity Invariance Detection Model (DIDM), which has two main components: the Diversity Learning Module (DLM) and the Weighted Aligning Module (WAM). These components work together to balance domain-invariant feature learning and feature diversity. Specifically, the DLM is designed to preserve feature diversity and restricting semantic information for domain-specific features, while the WAM aligns features across domains to enhance domain invariance.
3.1 Preliminaries
Data Augmented. For the single-domain generalization task, training a model on single source dataset may lead to overfitting to domain-specific styles and features. Besides, obtaining diverse datasets across multiple domains is often costly. To address this, we apply data augmentation techniques [1, 16, 24] to increase the diversity of the source domain images, thereby enhancing the model’s adaptability and robustness to variations across different domains. These augmentations include techniques such as noise addition, blurring, and random cropping. Each source image $x_{s}$ is transformed through these augmentation functions $A(.)$ to produce the enhanced image $x_{a}=A(x_{s})$.
3.2 Overview
@@
The net structure of the proposed model is shown in Fig. 2, where Faster R-CNN [33] is used as the base detector. Given a source image $x_{s}$, we can generate an argument image $x_{a}=A(x_{s})$. During the training phase, both $x_{s}$ and $x_{a}$ are fed into the model to produce their corresponding feature maps $F_{s}$ and $F_{a}$, respectively. Only the $F_{s}$ is fed into RPN to get a series of proposals $P\in R^{K\times 4}$. Then, the $F_{s}$, $F_{a}$ and $P$ are fed into the RoI-Pooling layer to get the features of proposals for the original and argument image: $z_{s}=RP(F_{s},P)$, $z_{a}=RP(F_{a},P)$.
@@
After that, the pooled feature $z_{s}$ and $z_{a}$ are fed into the further network for detection results, supervised by the loss function of Faster-RCNN. To support domain alignment and feature diversity learning, $z_{s}$ and $z_{a}$ are sent into the Weighted Aligning Module (WAM) and Diversity Learning Module (DLM). The goal of this process is to achieve domain alignment while preserving domain-specific diversity. In the DLM, feature diversity loss is applied to enhance the feature diversity, while entropy maximization loss eliminates the semantic information. In the WAM, the loss weight is used to prevent overemphasizing the domain alignment, which may influence the feature diversity. By integrating the supervision from both the DLM and WAM, the proposed Diversity Invariance Detection Model (DIDM) can achieve improved performance on the domain generalization task.
3.2.1 Diversity Learning Module (DLM)

### v2

@@ Abstract
Single-domain generalization for object detection (S-DGOD) seeks to transfer learned representations from a single source domain to unseen target domains. While recent approaches have primarily focused on achieving feature invariance, they ignore that domain diversity also presents significant challenges for the task. First, such invariance-driven strategies often lead to the loss of domain-specific information, resulting in incomplete feature representations. Second, cross-domain feature alignment forces the model to overlook domain-specific discrepancies, thereby increasing the complexity of the training process. To address these limitations, this paper proposes the Diversity Invariant Detection Model (DIDM), which achieves a harmonious integration of domain-specific diversity and domain invariance. Our key idea is to learn the invariant representations by keeping the inherent domain-specific features. Specifically, we introduce the Diversity Learning Module (DLM). This module limits the invariant semantics while explicitly enhancing domain-specific feature representation through a proposed feature diversity loss. Furthermore, to ensure cross-domain invariance without sacrificing diversity, we incorporate the Weighted Aligning Module (WAM) to enable feature alignment while maintaining the discriminative domain-specific information. Extensive experiments on multiple diverse datasets demonstrate the effectiveness of the proposed model, achieving superior performance compared to existing methods.
Index Terms:
Single-Domain Generalization, Object Detection, Transfer Learning
publicationid: pubid: 0000–0000/00\$00.00 © 2025 IEEE
I Introduction
@@
In recent years, deep learning has led to significant improvements in the performance of detection models [1, 2, 3, 4]. However, traditional detectors often rely on the assumption of Independent and Identically Distributed (i.i.d.) data. This assumption becomes a limitation when the model encounters a domain shift [5, 6], leading to a significant drop in performance and restricting the applicability of these models in real-world scenarios.
To address the domain shift challenge, domain adaptation (DA) methods were developed with the objective of improving model performance in the target domain by minimizing the domain discrepancy between source and target domains. However, the unavailability of the true label of the target domain has led researchers to indirectly manage target risk through source domain optimization while explicitly accounting for domain distributional differences [7]. According to the theoretical framework established by Ben-David et al. [8], DA typically employs feature alignment strategies to reduce the statistical divergence between domains. This approach, grounded in domain-invariant representation learning [9, 10, 11, 12], has become a central research direction in domain adaptation and domain generalization, simultaneously driving methodological innovations in these fields.
Building upon domain adaptation (DA) methods, existing domain generalization (DG) models also aim to reduce domain bias by learning domain-invariant features. By minimizing the risk across multiple source domains, these models extract features that are universally applicable to all source domains, thereby enhancing their generalization capability to unseen target domains. However, these approaches face three key limitations: information loss, conflicting constraints, and theoretical imbalances. Firstly, an overemphasis on invariance may hinder the model’s ability to capture domain-specific information, thereby compromising the completeness of feature representation. Secondly, this approach lacks rationality, as it requires the model to extract similar or even identical features from source domains that exhibit significant visual or semantic differences. Given the inherent diversity across domains, such constraints are counterintuitive and may hinder effective feature extraction, potentially complicating the training process. Finally, as illustrated in Fig. 1, the feature distributions of three scenarios are presented: (a) unaligned features, (b) fully aligned features without considering domain-specific characteristics, and (c) aligned features with domain-specific information preserved. Unlike traditional DA methods, where target domain samples are available for alignment, DG faces the challenge of unseen target domains. Simply aligning features from multiple source domains may overly constrain the feature distribution, leading to limited overlap with the probable target domain distribution. As shown in Fig. 1(c), the DG task must also account for domain-specific information to enable the model to achieve greater alignment with the target domain. To address these challenges, we propose that an effective DG model should synergistically optimize the learning of domain-invariant features while preserving and leveraging domain-specific diversity. Learning domain-invariant features ensures robustness across diverse domains, whereas fully exploiting domain-specific diversity significantly enhances feature representation and optimizes the training process. The ultimate goal is to construct a more inclusive and representative feature space that maximizes overlap with the target domain distribution.
Figure 1: Illustrate three scenarios of feature distribution: (a) unaligned features, (b) fully aligned features, (c) aligned features with domain-specific information preserved. The essence of feature alignment is to minimize the difference in distribution $\gamma$ between the source domain combination and the target domain. However, as illustrated in Fig. (b), forced alignment may result in an overly homogeneous feature space, neglecting the unique structural information inherent to each domain and failing to effectively cover the feature distribution of the target domain. Therefore, it is essential to consider the diversity of domain-specific information. As demonstrated in Fig. (c), this diversity-protecting alignment strategy can create a more inclusive and representative feature space, significantly enhancing the overlap between the distribution of combined features and the target domain.
@@
Based on the above discussion, specifically, in this paper, we introduce the Diversity Invariance Domain Generalization Detection Model (DIDM), designed to not only learn the domain invariance but also keep the diversity domain-specific. To manage these two feature types, we incorporate the Dive-
rsity Learning Module (DLM) to capture domain-specific features and the Weighted Aligning Module (WAM) to focus on domain-invariant features.
The Diversity Learning Module (DLM) is specifically designed to preserve the diversity of domain-specific features during the training phase. First, to enhance the distinctiveness of domain-specific features, the module explicitly separates domain-invariant information from domain-specific characteristics. Specifically, since domain-specific features should not contain common semantics across domains, we propose to suppress shared semantic representations by estimating domain-specific category semantics. This mechanism ensures that the learned features retain unique domain characteristics. Second, to further enhance the diversity of domain-specific features, we propose a feature diversity loss function. This loss explicitly encourages the model to capture and enrich the representational diversity of domain-specific features during training. By separating and enriching the domain-specific features, our model can learn more discriminative features, leading to an expanded and more robust feature distribution for the target domain.
In addition to preserving the diversity of domain-specific features, achieving domain-invariant representations is also critical for domain generalization (DG) tasks. However, overemphasizing domain-invariant features may also lead to the loss of discriminative domain-specific characteristics, thereby harming the overall feature diversity. To address this challenge, we introduce the Weighted Aligning Module (WAM) to achieve feature alignment while maintaining the inherent diversity of domain-specific features. Considering that excessive alignment often causes the collapse of feature diversity, we propose that features from different domains should retain a certain level of discrepancy. Specifically, we argue that if the features can achieve similar detection performance, further alignment may not be necessary. To this end, the WAM incorporates an adaptive loss weighting mechanism that dynamically balances the trade-off between alignment and diversity. This mechanism prevents the model from overemphasizing feature alignment and instead focuses on the most relevant features for alignment. By co-optimizing the Diversity Learning Module (DLM) and the WAM, our model achieves effective domain generalization on detection tasks. In conclusion, the contributions of the proposed Domain-Invariant Diversity Module (DIDM) are summarized as follows:
  • We propose the Diversity Learning Module (DLM) to preserve the inherent feature diversity within domain-specific representations. Additionally, this paper analyze the theoretical significance of feature diversity in domain generalization (DG) tasks through the transfer theory.
  • The Weighted Aligning Module (WAM) is designed to achieve cross-domain feature alignment while maintaining the intrinsic diversity of domain-specific representations.
  • Extensive experimental evaluations on diverse benchmark datasets demonstrate the robustness and generalization capability of our proposed method, validating its effectiveness across multiple scenarios.
II Related Work
Domain Generalization (DG). Researchers have been working extensively in single domain generalization and have developed numerous methods to enhance it [13, 14, 15, 16, 17]. These methods can generally be categorized into three main groups: data augmentation, presentation learning of domain invariance, and various learning strategies. First, the data argumentation approach [18, 19, 20] increases the diversity of source domain data by applying various data manipulation techniques. For instance, Somavarapu et al. [20] introduce a straightforward image stylization transform to generate diverse samples, exploring variability across sources. Similarly, Wang et al. [21] employ adversarial training to create varied input images, further enhancing generalization. Second, some kinds of DG methods focus on the presentation learning of domain invariance. Shao et al. [22] use multi-adversarial discriminative training to extract both shared and distinctive feature representations across multiple source domains. Last, Various learning strategies [23, 24, 25, 26, 27, 28] are employed to improve model generalization. For example, Zhao et al. [28] apply a meta-learning approach that simulates the model’s adaptation to new, unseen domains during training, thereby boosting its adaptability in unfamiliar environments. These approaches illustrate significant advancements in domain generalization, highlighting different techniques designed to improve models’ performance on previously unseen domains.
Object Detection. To enhance the model’s performance in object detection, researchers have conducted numerous experiments and proposed various methods in object detection. One prominent approach is Domain Adaptive Object Detection (DAOD) [29, 30, 31, 32, 33, 34], which enhances the model’s robustness in new domains by simultaneously training on both source and target domains. For instance, a common strategy involves minimizing the distance between global and local features, as outlined in [35, 5, 30, 36]. However, these methods require access to the target domain during model training, which imposes limitations on practical applications. Therefore, researchers have proposed Domain Generalized Object Detection (DGOD) [37, 38]. For instance, Chang et al. [37] enhance the model’s generalization ability by decoupling depth estimation from dynamic perspective enhancement. However, the effectiveness of DGOD largely depends on the number of accessible source domains, and collecting multiple source domains can be costly. Consequently, researchers have introduced the more challenging Single Domain Generalized Object Detection (S-DGOD) [39, 40, 41, 42, 43], which is generally categorized into feature normalization and invariant-based methods. Firstly, feature normalization methods such as IBN-Net, proposed by Pan et al. [42], enable the network to adjust its normalization strategy according to different tasks and datasets by combining Instance Normalization (IN) and Batch Normalization (BN). IterNorm, introduced by Huang et al. [44], avoids feature decomposition through a Newton iteration method, thereby improving the efficiency of the normalization process. Secondly, invariant-based approaches, such as UFR proposed by Liu et al. [14], enhance the model’s generalization performance by eliminating prototype bias and attentional bias. Wu et al. [45] propose a cyclic disentanglement self-distillation approach specifically for single-domain generalization in object detection, which enhances feature disentanglement. These approaches demonstrate significant progress in the field of target detection and contribute to the rapid advancement of the discipline.
III Methodology
In this section, we introduce the proposed Diversity Invariance Detection Model (DIDM), which comprises two key components: the Diversity Learning Module (DLM) and the Weighted Aligning Module (WAM). These components are synergistically designed to balance domain-invariant feature learning and feature diversity. Specifically, the DLM is engineered to preserve the intrinsic diversity of domain-specific features. Meanwhile, the WAM facilitates cross-domain feature alignment to promote domain invariance.
III-A Preliminaries
Theoretical Analysis. To address domain bias and optimize risk $\xi^{t}(h)$ in the target domain, domain adaptation typically aligns the probability distributions of the source $P_{X}^{s}$ and target $P_{X}^{t}$ domains while also constraining the risk $\xi^{s}(h)$ associated with the source domain. This mechanism is theoretically grounded in the Ben-David generalization bound [8, 7].
$\xi^{t}(h)\leq\xi^{s}(h)+d_{\mathcal{H}\Delta\mathcal{H}}(P_{X}^{s},P_{X}^{t})+\lambda_{\mathcal{H}}$ (1)
Where $\lambda_{\mathcal{H}}:=inf_{h\in\mathcal{H}}[\xi^{s}(h)+\xi^{t}(h)]$ indicates the ideal joint risk and $h$ represents an arbitrary classifier within the hypothesis space. $d_{\mathcal{H}\Delta\mathcal{H}}(P_{X}^{s},P_{X}^{t}):=sup_{h,h^{\prime}\in\mathcal{H}}|\xi^{s}(h,h^{\prime})-\xi^{t}(h,h^{\prime})|$ is utilized to measure inter-domain distributional differences. With domain adaptation theory, many domain adaptation methods aim to minimize the domain divergence $d_{\mathcal{H}\Delta\mathcal{H}}(P_{X}^{s},P_{X}^{t})$ through feature alignment, achieving great performance. These approaches have been increasingly adopted in domain generalization tasks in recent years, demonstrating their effectiveness in bridging distributional gaps across heterogeneous domains.
However, the theoretical foundation of domain generalization models differs from that of domain adaptation frameworks. While DA primarily focuses on reducing domain discrepancy between source and target domains, DG theory emphasizes the broader challenge of not only domain discrepancy but also domain diversity. Specifically, the theory domain generalization is presented as follows [46, 7]:
$\xi^{t}(h)\leq\sum\limits_{i=1}^{M}\pi_{i}^{*}\xi^{i}(h)+\frac{\gamma+\rho}{2}+\lambda_{H,(P_{X}^{t},P_{X}^{*})}$ (2)
$\gamma:=min_{\pi\in\Delta M}d_{\mathcal{H}}(P_{x}^{t},\sum\limits_{i=1}^{M}\pi_{i}P_{X}^{i})$ (3)
$\rho:=sup_{P_{X}^{\prime},P_{X}^{\prime\prime}\in\delta}d_{\mathcal{H}}(P_{X}^{\prime},P_{X}^{\prime\prime})$ (4)
$\delta:=\{\sum\limits_{i=1}^{M}\pi_{i}P_{X}^{i}|\pi\in\Delta_{M}\}$ (5)
Here, $M$ denotes the number of source domains, $P_{X}$ represents the feature distribution of a source domain, $\rho$ is the diameter of $\delta$, and $\gamma$ indicates the distance between $P_{x}^{t}$ and the joint feature distribution $\delta$, $P_{X}^{*}$ is the best approximating distribution.
The theory reveals that two key factors influence the target error bound in domain generation task: (1) the domain discrepancy among source domains ($\rho$), and (2) the distance between the joint feature distribution $\delta$ and the target domain ($\gamma$). Conventional methods primarily focus on minimizing the domain discrepancy $\rho$ by aligning source domains. However, this approach may inadvertently reduce the coverage of the joint feature distribution $\delta$, thereby increasing the distance between the joint distribution and the target domain. This trade-off can degrade generalization performance on unseen target domains and make the model to achieve suboptimal performance. To address this, we propose the DIDM framework, which simultaneously optimizes two objectives: (1) learning domain-invariant features to reduce $\rho$, and (2) enhancing the diversity of domain-specific information to expand the coverage of $\delta$. By balancing these goals, our method reduces both $\gamma$ and $\rho$, ultimately tightening the target error bound and improving model robustness to unseen domains.
@@ III-A Preliminaries
III-B Overview
@@
The net structure of the proposed model is shown in Fig. 2, where Faster R-CNN [1] is used as the base detector. Given a source image $x_{s}$, we can generate an argument image $x_{a}=A(x_{s})$. During the training phase, both $x_{s}$ and $x_{a}$ are fed into the model to produce their corresponding feature maps $F_{s}$ and $F_{a}$, respectively. Only the $F_{s}$ is fed into RPN to get a series of proposals $P\in R^{K\times 4}$. To suppress background influence, background boxes are filtered from the proposals $P$, retaining only the foreground boxes $P^{\prime}$. Then, the $F_{s}$, $F_{a}$ and $P^{\prime}$ are fed into the RoI-Pooling layer to get the features of proposals for the original and argument image: $z_{s}=RP(F_{s},P^{\prime})$, $z_{a}=RP(F_{a},P^{\prime})$.
@@
After the pooling layer, the pooled feature $z_{s}$ and $z_{a}$ are fed into the downstream detection network, where the detection results are supervised by the loss function of Faster-RCNN. To facilitate domain alignment and promote feature diversity, $z_{s}$ and $z_{a}$ are sent into the Weighted Aligning Module (WAM) and Diversity Learning Module (DLM). The goal of this design is to achieve domain alignment while preserving domain-specific diversity. In the DLM, a feature diversity loss is introduced to enhance the variation of domain-specific features, ensuring robustness to unseen domains. Simultaneously, an entropy maximization loss is applied to estimate the category semantic and ensure the domain-specific information. In the WAM, dynamic loss weighting is employed to balance the contributions of domain alignment and diversity preservation. By integrating the supervision from both the DLM and WAM, the Diversity Invariance Detection Model (DIDM) achieves improved performance on the domain generalization task.
III-C Diversity Learning Module (DLM)
Existing generalization methods typically focus on learning domain-invariant features, which often leads to the suppression or loss of domain-specific information, resulting in suboptimal performance. To address this limitation, the Diversity Learning Module (DLM) is introduced. First, domain-specific features are extracted from the features by employing a two-branch network structure to compare and extract domain-specific characteristics. To fully leverage domain-specific patterns, an entropy maximization loss is applied to avoid category semantics in the domain-specific features. Second, a diversity loss is introduced to enhance the variability of domain-specific features. This loss explicitly promotes feature diversity, ensuring that domain-specific representations remain distinct and robust. By combining the entropy maximization loss and diversity loss, the DLM effectively enhances the diversity of domain-specific information, leading to improved generalization performance in domain generalization detection tasks.